# Densities Of Primes And Primitive Roots

### v9

@@ 4.2 Finite Fourier Transforms and Equivalent Exponential Sums
since for each $d\mid p-1$, the map $t\longrightarrow z=r(d,t)\equiv dt\bmod p$ permutes the elements in $\mathbb{F}_{p}$. where the implied constant depends on the number of divisors of p1p-1. ∎
@@
$\sum_{n\in\mathcal{A}}e^{i2\pi b\tau^{n}/p}=\sum_{n\in\mathcal{A}}e^{i2\pi\tau^{n}/p}+O(p^{1/2+\delta}\log^{c}p),$ (42)
@@ 4.3 Finite Summation Kernels And Gaussian Sums
    $\sum_{n\leq p-1}\omega^{tn}=\frac{\omega^{t}-\omega^{tp}}{1-\omega^{t}},$
@@
    $\bigg{|}\sum_{n\leq p-1}\omega^{tn}\bigg{|}\leq\frac{2p}{\pi t}.$
@@
$\sum_{n\leq p-1}\omega^{tn}$ $=$ $\frac{\omega^{t}-\omega^{tp}}{1-\omega^{t}}.$
(ii) Observe that the parameters are $p$ prime, $\omega=e^{i2\pi/p}$, the integers $t\in[1,p-1]$, and $d\leq p-1<q-1$. This data implies that $\pi t/q\neq k\pi$ with $k\in\mathbb{Z}$, so the sine function $\sin(\pi t/p)\neq 0$ is well defined. Using standard manipulations, and $z/2\leq\sin(z)<z$ for $0<|z|<\pi/2$, the last expression becomes
$\bigg{|}\frac{\omega^{t}-\omega^{tp}}{1-\omega^{t}}\bigg{|}\leq\bigg{|}\frac{2}{\sin(\pi t/p)}\bigg{|}\leq\frac{2p}{\pi t}.$ (43)
@@
    $\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}\omega^{tn}=\sum_{d\mid p-1}\mu(d)\frac{\omega^{dt}-\omega^{dt((p-1)/d+1)}}{1-\omega^{dt}},$
@@
    $\left|\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}\omega^{tn}\right|\ll\frac{p^{1+\delta}}{\pi r(d,t)},$
    where $r(d,t)\equiv dt\bmod p$ is a permutation of $t\in\mathbb{F}_{p}$ and $\delta>0$ is a small number, for any $t\in[1,p-1]$.
@@
(ii) Observe that the parameters are $p$ prime, $\omega=e^{i2\pi/p}$, the integers $t\in[1,p-1]$ and $d\leq p-1<p$. In addition, the map $t\longrightarrow dt$ is a permutation in the finite field $\mathbb{F}_{p}$ for any fixed $d\neq 0$. This data implies that $\pi dt/p\neq k\pi$ with $k\in\mathbb{Z}$, so the sine function $\sin(\pi dt/p)\neq 0$ is well defined. Using standard manipulations, and $z/2\leq\sin(z)<z$ for $0<|z|<\pi/2$, the last expression becomes
$\left|\frac{\omega^{dt}-\omega^{dtp}}{1-\omega^{dt}}\right|\leq\left|\frac{2}{\sin(\pi dt/p)}\right|\leq\frac{2p}{\pi r(d,t)},$ (45)
@@
$\left|\sum_{d\mid p-1}\mu(d)\frac{\omega^{dt}-\omega^{dt((p-1)/d+1)}}{1-\omega^{dt}}\right|$ $\leq$ $\sum_{d\mid p-1}\left|\frac{2}{\sin(\pi dt/p)}\right|$ (46)
@@ 5 Evaluation Of The Main Term
([43, Lemma 1]) Let $x\geq 1$ be a large number, and let $\varphi(n)$ be the Euler totient function. Then
$\sum_{p\leq x}\frac{\varphi(p-1)}{p-1}=a_{1}\operatorname{li}(x)+O\left(\frac{x}{\log^{B}x}\right),$ (49)
@@
A more general version of this Lemma is proved in [45], and related discussions are given in [30, p. 16]. The generalization to number fields appears in [24]. These results are ubiquitous in various results in Number Theory. Here, the logarithm integral is defined by
$\operatorname{li}(x)=\int_{2}^{x}\frac{1}{\log t}dt=\frac{x}{\log x}+O\left(\frac{x}{\log^{2}x}\right).$ (50)
@@
$a_{1}=\prod_{p\geq 2}\left(1-\frac{1}{p(p-1)}\right)$ (51)
@@
$\sum_{p\leq x}\frac{1}{p}\sum_{\gcd(n,p-1)=1}1=a_{1}\operatorname{li}(x)+O\left(\frac{x}{\log^{B}x}\right).$ (52)
@@
$\sum_{p\leq x}\frac{\varphi(p-1)}{p-1}-\sum_{p\leq x}\frac{\varphi(p-1)}{p(p-1)}$ $=$ $a_{1}\operatorname{li}(x)+O\left(\frac{x}{\log^{B}x}\right)-\sum_{p\leq x}\frac{\varphi(p-1)}{p(p-1)}$ (54)
@@
$\sum_{p\leq x}\frac{\varphi(p-1)}{p(p-1)}=O\left(\log\log x\right)$ (55)
@@ 6 Estimate For The Error Term
$\sum_{x\leq p\leq 2x}\frac{1}{p}\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}\sum_{0<k\leq p-1}\psi\left((\tau^{n}-u)k\right)\ll\frac{x^{1-\varepsilon}}{\log x}$ (56)
@@
$\bigg{|}\widehat{V(u)}\bigg{|}$ $=$ $\left|\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{i2\pi\tau^{n}/p}+O\left(p^{1/2+\delta}(\log p)^{2}\right)\right|$
$\ll$ $\left|\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{i2\pi\tau^{n}/p}\right|+p^{1/2+\delta}(\log p)^{2}$
@@ 7 Densities Of Primes And Primitive Roots
$0=\sum_{x\leq p\leq 2x}\Psi(u).$ (61)
Replacing the characteristic function, Lemma 3.3, and expanding the nonexistence equation (61) yield
@@
where $B>1$. However, $\delta(u)>0$ contradicts the hypothesis (61) for all sufficiently large numbers $x\geq x_{0}$. Ergo, the short interval $[x,2x]$ contains primes with the fixed primitive root $u$. ∎
A formula for computing the density constant $\delta(u)\geq 0$, which is the density of primes with a fixed primitive root $u\neq\pm 1,v^{2}$, is derived in [23, p. 220]. This formula specifies the density as follows. Let $u=(st^{2})^{k}\neq\pm 1,v^{2}$ with $s$ squarefree, and $k\geq 1$, and let
$a_{k}(u)=\prod_{p\mid k}\frac{1}{p-1}\prod_{p\nmid k}\left(1-\frac{1}{p(p-1)}\right).$ (64)
@@
    $\delta(u)=\left(1-\mu(s)\prod_{\begin{subarray}{c}p\mid s\\ p\mid k\end{subarray}}\frac{1}{p-2}\prod_{\begin{subarray}{c}p\mid s\\ p\nmid k\end{subarray}}\frac{1}{p^{2}-p-1}\right)a_{k}(u).$ (66)
In the case of a decimal base $u=10$, the density (51), which is computed in [47], is
$\delta(10)=\prod_{p\geq 2}\left(1-\frac{1}{p(p-1)}\right)=0.37739558136192022880547280\ldots.$ (67)
@@
Information on the determination of the constant $\delta(u)\geq 0$, which is the density of primes with a fixed primitive root $u\neq\pm 1,v^{2}$, is provided in the next section and in [23, p. 220].
@@ 8 Application To Repeated Decimals
$\pi_{10}(x)=\#\left\{p\leq x:\operatorname{ord}_{p}(10)=p-1\right\}=\delta(10)\frac{x}{\log x}+O\left(\frac{x}{\log^{2}x}\right),$ (69)
where $\delta(10)=0.3773955\ldots$ is a constant, see (67), for all large numbers $x\geq 1$.
@@
$\frac{1}{p}=\frac{m}{10^{d}}+\frac{m}{10^{2d}}+\cdots=m\sum_{n\geq 1}\frac{1}{10^{dn}}=\frac{m}{10^{d}-1}$ (70)
@@
$10^{p-1}-1\equiv 0\bmod p^{2},$ (71)
are also important in the calculation of the period modulo $p^{2}$, confer [39, p. 333], [11] for other details.
@@
$\pi_{10}(x)=\sum_{p\leq x}\Psi(10)=\delta(10)\operatorname{li}(x)+O\left(\frac{x}{\log^{B}x}\right),$ (72)

### v10

@@ 4.2 Finite Fourier Transforms and Equivalent Exponential Sums
where the implied constant depends on the number of divisors of p1p-1. ∎
@@
$\sum_{n\in\mathcal{A}}e^{i2\pi b\tau^{n}/p}=\sum_{n\in\mathcal{A}}e^{i2\pi\tau^{n}/p}+O(p^{1/2+\delta}\log^{c}p),$ (4.27)
@@ 4.3 Finite Summation Kernels And Gaussian Sums
    $\bigg{|}\sum_{n\leq p-1}\omega^{tn}\bigg{|}\leq\frac{2p}{\pi t}.$
@@
    $\sum_{1\leq t\leq p-1}\bigg{|}\sum_{n\leq p-1}\omega^{tn}\bigg{|}\leq 2p\log p.$
@@
$\sum_{n\leq p-1}\omega^{tn}$ $=$ $\frac{\omega^{t}-\omega^{tp}}{1-\omega^{t}}\leq\bigg{|}\frac{\omega^{t}-\omega^{tp}}{1-\omega^{t}}\bigg{|}\leq\bigg{|}\frac{2}{\sin(\pi t/p)}\bigg{|}\leq\frac{2p}{\pi t}.$ (4.28)
This follows from $t\in[1,p-1]$, the sine function $\sin(\pi t/p)\neq 0$ is well defined and $z/2\leq\sin(z)<z$ for $0<|z|<\pi/2$.
(ii) Using last expression (i) this becomes
$\sum_{1\leq t\leq p-1}\bigg{|}\sum_{n\leq p-1}\omega^{tn}\bigg{|}$ $\leq$ $\sum_{1\leq t\leq p-1}\frac{2p}{\pi t}\leq 2p\log p.$ (4.29)
@@
    $\Bigg{|}\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}\omega^{tn}\Bigg{|}\ll\frac{p^{1+\delta}}{\pi r(d,t)},$
    where $r(d,t)\equiv dt\bmod p$ is a permutation of $t\in\mathbb{F}_{p}$.
@@
    $\sum_{1\leq t\leq p-1}\Bigg{|}\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}\omega^{tn}\Bigg{|}\ll p^{1+\delta}\log p,$
    where $\delta>0$ is a small number.
@@
Now, the parameters are $p$ prime, $\omega=e^{i2\pi/p}$, the integers $t\in[1,p-1]$ and $d\leq p-1<p$. In addition, the map $t\longrightarrow dt$ is a permutation in the finite field $\mathbb{F}_{p}$ for any fixed $d\neq 0$. This data implies that $\pi dt/p\neq k\pi$ with $k\in\mathbb{Z}$, so the sine function $\sin(\pi dt/p)\neq 0$ is well defined. Using standard manipulations, and $z/2\leq\sin(z)<z$ for $0<|z|<\pi/2$, the last expression becomes
$\left|\frac{\omega^{dt}-\omega^{dtp}}{1-\omega^{dt}}\right|\leq\left|\frac{2}{\sin(\pi dt/p)}\right|\leq\frac{2p}{\pi r(d,t)},$ (4.31)
@@
$\left|\sum_{d\mid p-1}\mu(d)\frac{\omega^{dt}-\omega^{dt((p-1)/d+1)}}{1-\omega^{dt}}\right|$ $\leq$ $\sum_{d\mid p-1}\left|\frac{2}{\sin(\pi dt/p)}\right|$
@@
where $\sum_{d\mid p-1}1=d(p-1)\ll p^{\delta}$ is the number of divisor in p1p-1 and $\delta>0$ is a small number.
(ii) For each $d\mid p-1$, the map $t\longrightarrow z\equiv dt\bmod p$ is a permutation in the finite field $\mathbb{F}_{p}$. Thus,
$\sum_{1\leq t\leq p-1}\Bigg{|}\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}\omega^{tn}\Bigg{|}$ $\leq$ $\sum_{d\mid p-1,}\sum_{1\leq t\leq p-1}\left|\frac{2}{\sin(\pi dt/p)}\right|$
$\leq$ $\sum_{d\mid p-1,}\sum_{1\leq z\leq p-1}\frac{2p}{\pi z}$
$\ll$ $p^{1+\delta}\log p,$
@@ 5 Evaluation Of The Main Term
([42, Lemma 1]) Let $x\geq 1$ be a large number, and let $\varphi(n)$ be the Euler totient function. Then
$\sum_{p\leq x}\frac{\varphi(p-1)}{p-1}=a_{1}\operatorname{li}(x)+O\left(\frac{x}{\log^{B}x}\right),$ (5.1)
@@
A more general version of this Lemma is proved in [44], and related discussions are given in [29, p. 16]. The generalization to number fields appears in [23]. These results are ubiquitous in various results in Number Theory. Here, the logarithm integral is defined by
$\operatorname{li}(x)=\int_{2}^{x}\frac{1}{\log t}dt=\frac{x}{\log x}+O\left(\frac{x}{\log^{2}x}\right).$ (5.2)
@@
$a_{1}=\prod_{p\geq 2}\left(1-\frac{1}{p(p-1)}\right)$ (5.3)
@@
$\sum_{p\leq x}\frac{1}{p}\sum_{\gcd(n,p-1)=1}1=a_{1}\operatorname{li}(x)+O\left(\frac{x}{\log^{B}x}\right).$ (5.4)
@@
$\sum_{p\leq x}\frac{\varphi(p-1)}{p-1}-\sum_{p\leq x}\frac{\varphi(p-1)}{p(p-1)}$ $=$ $a_{1}\operatorname{li}(x)+O\left(\frac{x}{\log^{B}x}\right)-\sum_{p\leq x}\frac{\varphi(p-1)}{p(p-1)}$ (5.6)
@@
$\sum_{p\leq x}\frac{\varphi(p-1)}{p(p-1)}=O\left(\log\log x\right)$ (5.7)
@@ 6 Estimate For The Error Term
$\sum_{x\leq p\leq 2x}\frac{1}{p}\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}\sum_{0<k\leq p-1}\psi\left((\tau^{n}-u)k\right)\ll\frac{x^{1-\varepsilon}}{\log x}$ (6.1)
@@
$\bigg{|}\widehat{V(1,u)}\bigg{|}$ $=$ $\left|-\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{i2\pi\frac{\tau^{n}}{p}}+O(p^{1/2+\delta}\log^{2}p)\right|$
$\ll$ $\left|\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{i2\pi\frac{\tau^{n}}{p}}\right|+p^{1/2+\delta}(\log p)^{2}$
@@ 7 Densities Of Primes And Primitive Roots
$0=\sum_{x\leq p\leq 2x}\Psi(u).$ (7.2)
Replacing the characteristic function, Lemma 3.3, and expanding the nonexistence equation (7.2) yield
@@
where $B>1$. However, $\delta(u)>0$ contradicts the hypothesis (7.2) for all sufficiently large numbers $x\geq x_{0}$. Ergo, the short interval $[x,2x]$ contains primes with the fixed primitive root $u$. ∎
A formula for computing the density constant $\delta(u)\geq 0$, which is the density of primes with a fixed primitive root $u\neq\pm 1,v^{2}$, is derived in [22, p. 220]. This formula specifies the density as follows. Let $u=(st^{2})^{k}\neq\pm 1,v^{2}$ with $s$ squarefree, and $k\geq 1$, and let
$a_{k}(u)=\prod_{p\mid k}\frac{1}{p-1}\prod_{p\nmid k}\left(1-\frac{1}{p(p-1)}\right).$ (7.5)
@@
    $\delta(u)=\left(1-\mu(s)\prod_{\begin{subarray}{c}p\mid s\\ p\mid k\end{subarray}}\frac{1}{p-2}\prod_{\begin{subarray}{c}p\mid s\\ p\nmid k\end{subarray}}\frac{1}{p^{2}-p-1}\right)a_{k}(u).$ (7.7)
In the case of a decimal base $u=10$, the density (5.3), which is computed in [46], is
$\delta(10)=\prod_{p\geq 2}\left(1-\frac{1}{p(p-1)}\right)=0.37739558136192022880547280\ldots.$ (7.8)
@@
Information on the determination of the constant $\delta(u)\geq 0$, which is the density of primes with a fixed primitive root $u\neq\pm 1,v^{2}$, is provided in the next section and in [22, p. 220].
@@ 8 Application To Repeated Decimals
$\pi_{10}(x)=\#\left\{p\leq x:\operatorname{ord}_{p}(10)=p-1\right\}=\delta(10)\frac{x}{\log x}+O\left(\frac{x}{\log^{2}x}\right),$ (8.2)
where $\delta(10)=0.3773955\ldots$ is a constant, see (7.8), for all large numbers $x\geq 1$.
@@
$\frac{1}{p}=\frac{m}{10^{d}}+\frac{m}{10^{2d}}+\cdots=m\sum_{n\geq 1}\frac{1}{10^{dn}}=\frac{m}{10^{d}-1}$ (8.3)
@@
$10^{p-1}-1\equiv 0\bmod p^{2},$ (8.4)
are also important in the calculation of the period modulo $p^{2}$, confer [38, p. 333], [10] for other details.
@@
$\pi_{10}(x)=\sum_{p\leq x}\Psi(10)=\delta(10)\operatorname{li}(x)+O\left(\frac{x}{\log^{B}x}\right),$ (8.5)